OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!!!! The length of a rectangle is 7 m less than triple the width. If w represents the width, write an expression for the length, l.

OpenStudy (anonymous):

@aahmed123 @DestinyRocks!

OpenStudy (anonymous):

do you have the answer choices?

OpenStudy (anonymous):

Here are the answer choices: A. l = 1/3 w - 7 B. l = 3w – 7 C. l = 1/3 w + 7 D. l = 3w + 7

OpenStudy (anonymous):

ok so accordnig to the word problem, there is no addition

OpenStudy (anonymous):

and there are no fractions, so that leaves you with one answer

OpenStudy (anonymous):

That would make the answer B.

OpenStudy (anonymous):

yup :D

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no problem